# BlurDM: A Blur Diffusion Model for Image Deblurring

### v2

@@ A.2 ELBO and Optimization for BlurDM
To reconstruct the sharp image $I_{0}$ from the degraded observation $I_{T}$, we adopt the variational inference framework of DDPM [NEURIPS2020_4c5bcfec] and derive an evidence lower bound (ELBO) that explicitly incorporates the blur residuals $e_{1:T}$. The joint ELBO is expressed as
@@
Given this structural alignment between the forward marginal and the prior, we follow the standard approach of DDPM [NEURIPS2020_4c5bcfec] and DDIM [song2021denoising], and omit the terminal KL divergence $D_{\mathrm{KL}}(q(I_{T}\mid I_{0},e_{1:T})\parallel p_{\theta}(I_{T}))$ during training. Instead, we retain only the stepwise KL divergence terms, which capture the discrepancy between the reverse and forward transitions at each timestep as:
@@
In our setting, since the variances of the two Gaussian distributions are matched exactly, the KL divergence reduces to a squared difference between their means, as is standard in DDPM [NEURIPS2020_4c5bcfec]. Accordingly, the KL divergence term in (25) reduces to
@@
Although per‑step ground‑truth blur residuals are unavailable, recent diffusion‑based research [chen2023hierarchical, geng2025consistency, salimansprogressive, xia2023diffir] has shown that supervising the generated results is sufficient to train the diffusion model.
@@ A.3 Deterministic Implicit Sampling for BlurDM
We follow the approach of DDIM [song2021denoising] and proceed by mathematical induction from $t=T$ to $t=1$. Assuming that the marginal distribution $q(I_{t}\mid I_{0},e_{1:t})$ is valid at step $t$, we aim to prove that sampling $I_{t-1}$ from $q_{\sigma}(I_{t-1}\mid I_{t},I_{0}^{\theta},e_{1:t})$ yields a distribution consistent with
@@ A.7 Visualizations of Blur Residual
In Fig. 9, we compare blurred images at different exposure times synthesized from the GoPro [Nah_2017_CVPR] dataset and the corresponding blur residuals obtained using BlurDM. Here, we depict $T=0$ as sharp images. As $T$ increases (i.e., as the exposure time increases), the images gradually become more blurred, and the corresponding blur residuals evolve accordingly. Eventually, at $T=1$, the images correspond to the fully blurred versions.
@@
We visualize and compare blur residuals generated by BlurDM and RDDM [liu2024residual] in Fig. 8. RDDM computes blur residuals using a simple subtraction operation followed by linear scaling to adjust intensity, resulting in residuals that primarily capture direct differences between blurred and sharp images, progressively magnified over time steps ($e_{t}$ to $e_{t+2}$). In contrast, BlurDM addresses the inductive bias of blur formation process to estimate blur residuals, capturing the non-linear progressive accumulation of blur. Unlike RDDM, BlurDM models how blurred residuals spatially diffuse and evolve over time, simulating the blur spread as exposure time extends. As a result, BlurDM more accurately represents the physical characteristics of motion blur within the diffusion process, leading to significantly improved deblurring performance.
@@ A.9 Deblurred Results on Real-world Datasets
We provide additional deblurred results for deblurring models using BlurDM, compared to those without using our method, referred to as Baseline. These deblurring models are trained on the GoPro [Nah_2017_CVPR] and RealBlur-J [rim_2020_ECCV] training sets, and tested on the RealBlur-J testing set. We demonstrate qualitative comparisons based on four image deblurring models, including MIMO-UNet [MIMO] in Fig. 10, Stripformer [Tsai2022Stripformer] in Fig. 11, FFTformer [Kong_2023_CVPR] in Fig. 12, and LoFormer [mao2024loformer] in Fig. 13.
@@ A.10 Broader Impacts
Figure 9: Visualizations of blur residual from GoPro [Nah_2017_CVPR] dataset
Figure 10: Qualitative results of MIMO-UNet [MIMO] on the RealBlur-J [rim_2020_ECCV] dataset.
Figure 11: Qualitative results of Stripformer [Tsai2022Stripformer] on the RealBlur-J [rim_2020_ECCV] dataset.
Figure 12: Qualitative results of FFTformer [Kong_2023_CVPR] on the RealBlur-J [rim_2020_ECCV] dataset.
Figure 13: Qualitative results of LoFormer [mao2024loformer] on the RealBlur-J [rim_2020_ECCV] dataset.

### v3

@@ A.2 ELBO and Optimization for BlurDM
To reconstruct the sharp image $I_{0}$ from the degraded observation $I_{T}$, we adopt the variational inference framework of DDPM [10] and derive an evidence lower bound (ELBO) that explicitly incorporates the blur residuals $e_{1:T}$. The joint ELBO is expressed as
@@
Given this structural alignment between the forward marginal and the prior, we follow the standard approach of DDPM [10] and DDIM [38], and omit the terminal KL divergence $D_{\mathrm{KL}}(q(I_{T}\mid I_{0},e_{1:T})\parallel p_{\theta}(I_{T}))$ during training. Instead, we retain only the stepwise KL divergence terms, which capture the discrepancy between the reverse and forward transitions at each timestep as:
@@
In our setting, since the variances of the two Gaussian distributions are matched exactly, the KL divergence reduces to a squared difference between their means, as is standard in DDPM [10]. Accordingly, the KL divergence term in (25) reduces to
@@
Although per‑step ground‑truth blur residuals are unavailable, recent diffusion‑based research [4, 8, 35, 47] has shown that supervising the generated results is sufficient to train the diffusion model.
@@ A.3 Deterministic Implicit Sampling for BlurDM
We follow the approach of DDIM [38] and proceed by mathematical induction from $t=T$ to $t=1$. Assuming that the marginal distribution $q(I_{t}\mid I_{0},e_{1:t})$ is valid at step $t$, we aim to prove that sampling $I_{t-1}$ from $q_{\sigma}(I_{t-1}\mid I_{t},I_{0}^{\theta},e_{1:t})$ yields a distribution consistent with
@@ A.7 Visualizations of Blur Residual
In Fig. 9, we compare blurred images at different exposure times synthesized from the GoPro [24] dataset and the corresponding blur residuals obtained using BlurDM. Here, we depict $T=0$ as sharp images. As $T$ increases (i.e., as the exposure time increases), the images gradually become more blurred, and the corresponding blur residuals evolve accordingly. Eventually, at $T=1$, the images correspond to the fully blurred versions.
@@
We visualize and compare blur residuals generated by BlurDM and RDDM [18] in Fig. 8. RDDM computes blur residuals using a simple subtraction operation followed by linear scaling to adjust intensity, resulting in residuals that primarily capture direct differences between blurred and sharp images, progressively magnified over time steps ($e_{t}$ to $e_{t+2}$). In contrast, BlurDM addresses the inductive bias of blur formation process to estimate blur residuals, capturing the non-linear progressive accumulation of blur. Unlike RDDM, BlurDM models how blurred residuals spatially diffuse and evolve over time, simulating the blur spread as exposure time extends. As a result, BlurDM more accurately represents the physical characteristics of motion blur within the diffusion process, leading to significantly improved deblurring performance.
@@ A.9 Deblurred Results on Real-world Datasets
We provide additional deblurred results for deblurring models using BlurDM, compared to those without using our method, referred to as Baseline. These deblurring models are trained on the GoPro [24] and RealBlur-J [31] training sets, and tested on the RealBlur-J testing set. We demonstrate qualitative comparisons based on four image deblurring models, including MIMO-UNet [5] in Fig. 10, Stripformer [40] in Fig. 11, FFTformer [13] in Fig. 12, and LoFormer [21] in Fig. 13.
@@ A.10 Broader Impacts
Figure 9: Visualizations of blur residual from GoPro [24] dataset
Figure 10: Qualitative results of MIMO-UNet [5] on the RealBlur-J [31] dataset.
Figure 11: Qualitative results of Stripformer [40] on the RealBlur-J [31] dataset.
Figure 12: Qualitative results of FFTformer [13] on the RealBlur-J [31] dataset.
Figure 13: Qualitative results of LoFormer [21] on the RealBlur-J [31] dataset.